OpenStudy (anonymous):

multiple choice, Write an equation in point slope form for the line that passes through (-4, 3) and whose slope is 3 1)y – 3 = 3(x – 4) 2)y – 3 = 3(x + 4) 3)y – 4 = -3(x -3) 4)y + 4 = 3(x + 3)

OpenStudy (saifoo.khan):

3 = 3(-4) + c c = 3 + 12 c = 15

OpenStudy (saifoo.khan):

y = 3x + 15

OpenStudy (anonymous):

im confused, isn't the answer one of the multiple choice?

OpenStudy (saifoo.khan):

Solve them

OpenStudy (saifoo.khan):

2)y – 3 = 3(x + 4)

OpenStudy (anonymous):

i really don't know how man.

OpenStudy (saifoo.khan):

y = 3x + 12+3 y = 3x + 15

OpenStudy (saifoo.khan):

Lol, y dont u?? havnt u learn them yet?? then how u got this HW?

OpenStudy (anonymous):

This is the equation for point-slope form: \[y-y_1 = m(x-x_1)\] where \[(x_1,y_1)\] is the point given, and m is the slope. Since we are given the point (-4,3), and the slope is 3, we have: \[x_1 = -4, y_1 = 3, m = 3\] putting those into the equation we obtain: \[y-y_1=m(x-x_1) \Rightarrow y-3=3(x-(-4))\Rightarrow y-3=3(x+4)\] which i believe is answer choice 2.

OpenStudy (anonymous):

if you need anything explained in further detail just ask.